# Faster and Simpler Online Computation of String Net Frequency

### v2

@@ 3.4 Dealing with Type (i) node
Let $u$ be the parent of the leaf representing $S$ in $\mathsf{STree}(S)$. If $u\neq v$, then $u$ is a node of Type (i) and we have $(a,b)\in\Phi_{aS}(u)\setminus\Phi_{S}(u)$, where $b$ is the first character of the edge label from $u$ to $S$ (see also Figure 1). In addition, $u$ is the only possible node of Type (i).
@@
It follows from the definition of Type (i) nodes that for any Type (i) nodes, the first character $\alpha$ in the pair $(\alpha,\beta)$ has to be the new character $a$ that is prepended to $S$. If $(a,\beta)\in\Phi_{aS}(u)$ for the parent $u$ of the leaf $S$, then this character $\beta$ must be unique by Lemma 3. Further by Lemma 6, this $\beta$ is the first character $b$ of the edge label from $u$ to $S$ (see also the right diagram of Figure 1). Now consider the case that $u\neq v$, where $v$ is the lowest ancestor of $S$ that has a (soft or hard) Weiner link labeled $a$ in $\mathsf{STree}(S)$ before the update. Since $u\neq v$, $u$ obtains a new soft Weiner link ${\mathsf{W\_link}}_{a}({u})$ labeled $a$ toward the leaf $aS$ in $\mathsf{STree}(aS)$ after the update (see the right diagram of Figure 1). Then the conditions of Lemma 6 are satisfied for the node $u$ and the character pair $(a,b)$ in $\mathsf{STree}(aS)$. Therefore, $(a,b)\in\Phi_{aS}(u)$ holds in this case. This can also be observed by the corresponding new substructure in form of Figure 2 that appears in $\mathsf{STree}(aS)$. Clearly $(a,b)\notin\Phi_{S}(u)$ because $\mathsf{occ}_{S}(aub)=0$. Hence $u$ is of Type (i) when the aforementioned conditions are satisfied.
Next, we show that $u$ is the only possible node of Type (i). Let $u^{\prime}$ be any node on the path between $v$ and the leaf for $S$, exclusive. Notice that these nodes $u^{\prime}$ are the only nodes other than $u$ that obtain a new left-extension $a$ (see also Figure 1). Assume for a contrary that $u^{\prime}$ is of Type (i), which implies that $\mathsf{occ}_{aS}(au^{\prime})=1$ and $\mathsf{occ}_{S}(au^{\prime})=0$. Let $b^{\prime}$ be the only character that immediately follows $au^{\prime}$ in $aS$, namely $\mathsf{occ}_{aS}(au^{\prime}b^{\prime})=1$. Note that such character $b^{\prime}$ must exist for each $u^{\prime}$. It follows from Lemma 6 that it suffices to check whether $u^{\prime}$ has an out-edge labeled $b^{\prime}$ that leads to a leaf. However, this is not the case: $au^{\prime}b^{\prime}$ is a prefix of $aS$, and thus $u^{\prime}b^{\prime}$ is a prefix of $S$. Thus, $b^{\prime}$ is the first character in the path label from $u^{\prime}$ to the leaf $S$. Since $u^{\prime}$ is a non-parent ancestor of the leaf $S$, the out-edge from $u^{\prime}$ that begins with $b^{\prime}$ does not lead to a leaf. Thus, $u^{\prime}$ cannot be a Type (i) node by Lemma 6. ∎
This node $u$ is found in the process of climbing up the path from the leaf for $S$ toward its lowest ancestor $v$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({v})$ for $a\in\Sigma$. The character $b$ can easily be obtained by using the string depths of $u$ and $S$. When all the conditions in Lemma 7 are satisfied, then $u$ is of Type (i). We then update $\Phi_{aS}(u)\leftarrow\Phi_{S}(u)\cup\{(a,b)\}$ and replace $(u,\phi_{S}(u))$ with $(u,\phi_{S}(u)+1)=(u,\phi_{aS}(u))$ in the list $\mathcal{L}$.
@@ 3.5 Dealing with Type (ii) node
We begin with a characterization of a node $w$ of Type (ii):
Lemma 8.
For any node $w$ of Type (ii), $w\notin\mathsf{M}(S)$ and $w\in\mathsf{M}(aS)$ hold.
Proof.
By Case (ii) in Section 2.3, we have $\phi_{aS}(w)\geq 1$. By Lemma 1, $w\in\mathsf{M}(aS)$ holds.
Suppose for a contrary that $w\in\mathsf{M}(S)$. Then, since $\phi_{S}(w)=0$ (by Case (ii) in Section 2.3), for any pair $(\alpha,\beta)\in\Sigma\times\Sigma$ of characters we have $\mathsf{occ}_{S}(\alpha w)>1$ or $\mathsf{occ}_{S}(w\beta)>1$. By monotonicity, $\mathsf{occ}_{aS}(p)\geq\mathsf{occ}_{S}(p)$ holds for any strings $p,S$ and character $a$. Therefore, we have $\mathsf{occ}_{aS}(\alpha w)>1$ and $\mathsf{occ}_{aS}(w\beta)>1$ for any pair $(\alpha,\beta)$ of characters. However, this contradicts that $w$ has a positive NF in $aS$. ∎
@@
Figure 3: A snapshot of Weiner’s suffix tree construction algorithm in the case where $ax$ is a leaf. If the leaf $x$ is a child of $v$, then $(a,c)\in\Phi_{S}(v)\setminus\Phi_{aS}(v)$. Let $d$ be a unique character that immediately precedes $ax$ in $S$. Then $(d,c)\in\Phi_{aS}(av)\setminus\Phi_{S}(av)$. If $daz$ in an implicit node on the edge leading to the leaf $dax$, then $(d,e)\in\Phi_{S}(az)\setminus\Phi_{aS}(az)$.
@@
Recall that $av$ becomes an internal node as the parent of a new leaf for $aS$ in $\mathsf{STree}(aS)$ after the update. Since Weiner’s algorithm inserts at most one internal node and it is $av$ in this case, the string $dav$ remains on an edge in $\mathsf{STree}(aS)$ after the update. Thus, $av$ obtains a soft Weiner link labeled $d\in\Sigma$ toward node $dax$. Since $ax$ is a leaf, $dax$ is also a leaf. For the string $t$ such that $dax=davt$, we have $t[1]=c$. Thus, we have $(d,c)\in\Phi_{aS}(av)$ by Lemma 6. This can also be observed by the corresponding new substructure in form of Figure 2 that appears in $\mathsf{STree}(aS)$. Due to Lemma 3, $d$ is the unique character pair for $c$ that can possibly be a member of $\Phi_{aS}(av)$. Since $c$ was the only character immediately following $av$ in $S$, $(d,c)$ is the only character pair that could be a member for $\Phi_{aS}(av)$. Since $\mathsf{occ}_{S}(av)=1$, $av$ was not a repeat in $S$ before the update. Thus $\phi_{S}(av)=0$ and hence $(d,c)\notin\Phi_{S}(av)$ holds for the string $S$ before the update.
If $ax$ is not a leaf, then $\mathsf{occ}_{aS}(avc)\geq\mathsf{occ}_{aS}(ax)>1$, implying $(d,c)\notin\Phi_{aS}(av)$. Also, if $av$ is a node in $\mathsf{STree}(S)$, then $av$ was already a maximal repeat in $S$ and thus $av$ is not a Type (ii) node. ∎
The locus for $av$ is found via the Weiner link ${\mathsf{W\_link}}_{a}({v})$ from the node $v$ during the update from $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$. We remark that $av$ is on an edge in $\mathsf{STree}(S)$ and the node $ax$ is a leaf, if and only if ${\mathsf{W\_link}}_{a}({v})$ in $\mathsf{STree}(S)$ before the update is a soft Weiner link. We can easily check if $ax$ is a leaf or not. The character $c$ can be obtained by using the string depths of the nodes $av$ and $ax$ (see also Figure 3). The character $d$ can be found easily (e.g. by using the unique Weiner link from leaf $ax$.)
@@ 3.6 Dealing with Type (iii)-1 node
Consider the case that $\mathsf{occ}_{S}(av)=1$ and $\mathsf{occ}_{aS}(av)>1$. Since $\mathsf{occ}_{S}(av)=1$, we have $\mathsf{occ}_{S}(ax)=1$ and thus $ax$ is a leaf. Since $ax\in\mathsf{Suffix}(S)$, $x\in\mathsf{Suffix}(S)$ and hence $x$ terminates with \$currency-dollar\\$\$. Since $v$ is an internal node, $v$ cannot end with \$currency-dollar\\$\$ and thus $v$ is a proper prefix of $x$. Thus $av$ is on the edge leading to the leaf $ax$ in $\mathsf{STree}(S)$ before the update. Since $x$ terminates with \$currency-dollar\\$\$, $x$ is also a leaf. Let $c$ be the first character of the path label from $v$ toward $x$. If $x$ is not a child of $v$, then $\mathsf{occ}_{S}(vc)>1$ and thus $v$ cannot be of Type (iii)-1. Consider the case where $x$ is a child of $v$. In this case, $\mathsf{occ}_{S}(av)=\mathsf{occ}_{S}(vc)=\mathsf{occ}_{S}(avc)=1$ (see also the left diagram of Figure 3), and thus we have $(a,c)\in\Phi_{S}(v)$. After the update, $av$ becomes a branching node in $\mathsf{STree}(aS)$, meaning that $\mathsf{occ}_{aS}(av)>1$ (see also the right diagram of Figure 3). By Lemma 6, $(a,c)\notin\Phi_{aS}(v)$ holds for node $v$. This can also be observed with the corresponding existing substructure in form of Figure 2 that disappears in $\mathsf{STree}(aS)$. Overall, if the aforementioned conditions are satisfied, then $v$ is a node of Type (iii)-1. Since $\mathsf{occ}_{S}(avc)=1$, $c$ is the only right-extension of $av$ in $S$. It now follows from Lemma 3 that $a$ is the only pairing character for $c$, and hence $(a,c)$ is the only possible character pair such that $(a,c)\in\Phi_{S}(v)\setminus\Phi_{aS}(v)$.
Next, we show that $v$ is the only possible node that is of Type (iii)-1. Let $v^{\prime}$ be any node in the path between $v$ and $z$ exclusive, where $z$ is the lowest ancestor of the leaf $S$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ labeled $a$ in $\mathsf{STree}(S)$ before the update. Note that those nodes $v^{\prime}$ are the only nodes other than $v$ such that $\mathsf{occ}_{S}(av^{\prime})=1$ and $\mathsf{occ}_{S}(v^{\prime})>1$. However, since their descendant $v$ is an internal node, we have $\mathsf{occ}_{S}(v^{\prime}c^{\prime})>1$ and thus $(a,c^{\prime})\notin\Phi_{aS}(v^{\prime})$ holds, where $c^{\prime}$ is the only possible paring character for $a$ with node $v^{\prime}$ due to Lemma 3. Thus $v$ is the only possible node of Type (iii)-1. ∎
@@
Recall that $z$ is the lowest ancestor of the leaf $S$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ labeled $a$ in $\mathsf{STree}(S)$ before the update. If $v\neq z$, then $av$ is on an edge. We can easily check whether $x$ and $ax$ are leaves and $x$ is a child of $v$ in $\mathsf{STree}(S)$. Notice that the first condition $v\neq z$ is critical, since if $v=z$, then $av$ ($=az$) is an internal node in $\mathsf{STree}(S)$ and thus $\mathsf{occ}_{S}(av)>1$. Now we have two sub-cases:
@@ 3.7 Dealing with Type (iii)-2 node
Recall that $az$ is the parent of $ax$ in $\mathsf{STree}(S)$ before the update. Consider the case that $\mathsf{occ}_{S}(aze)=1$, where $ax$ is a leaf in $\mathsf{STree}(S)$. Also, consider the case that $\mathsf{occ}_{S}(daz)=1$. Since $daze$ is a prefix of $dax$, $\mathsf{occ}_{S}(daze)\geq 1$. Moreover, since $\mathsf{occ}_{S}(aze)=1$, we have $\mathsf{occ}_{S}(daze)=1$. Therefore, $(d,e)\in\Phi_{S}(az)$ by Lemma 6. Consider the case where $av$ becomes a branching node in $\mathsf{STree}(aS)$ after the update. Then, $av$ becomes a child of $az$, meaning that $\mathsf{occ}_{aS}(aze)>1$. Thus $(d,e)\notin\Phi_{aS}(az)$ holds for $az$. Overall, if the aforementioned conditions are satisfied, then $az$ is a node of Type (iii)-2. This can also be observed with the corresponding existing substructure in form of Figure 2 that disappears in $\mathsf{STree}(aS)$.
@@ 3.8 Complexities
By maintaining the list $\mathcal{L}$ for $\mathsf{NF^{+}}(S)$ in an online manner, we obtain the following:

### v3

@@ 3.4 Dealing with Type (i) node
Let $v$ be the lowest ancestor of $S$ such that the Weiner link ${\mathsf{W\_link}}_{a}({v})$ with the new character $a$ is defined, and $u$ be the parent of the leaf representing $S$ in $\mathsf{STree}(S)$. If $u\neq v$, then $u$ is a node of Type (i) and we have $(a,b)\in\Phi_{aS}(u)\setminus\Phi_{S}(u)$, where $b$ is the first character of the edge label from $u$ to $S$ (see also Figure 1). In addition, $u$ is the only possible node of Type (i).
@@
It follows from the definition of Type (i) nodes that for any Type (i) nodes, the first character $\alpha$ in the pair $(\alpha,\beta)$ has to be the new character $a$ that is prepended to $S$. If $(a,\beta)\in\Phi_{aS}(u)$ for the parent $u$ of the leaf $S$, then this character $\beta$ must be unique by Lemma 3. Further by Lemma 6, this $\beta$ is the first character $b$ of the edge label from $u$ to $S$ (see also the right diagram of Figure 1). Now consider the case that $u\neq v$, where $v$ is the lowest ancestor of $S$ that has a (soft or hard) Weiner link labeled $a$ in $\mathsf{STree}(S)$ before the update. Since $\neq v, uu$ obtains a new soft Weiner link ${\mathsf{W\_link}}_{a}({u})$ labeled $a$ toward the leaf $aS$ in $\mathsf{STree}(aS)$ after the update (see the right diagram of Figure 1). Then the conditions of Lemma 6 are satisfied for the node $u$ and the character pair $(a,b)$ in $\mathsf{STree}(aS)$. Therefore, $(a,b)\in\Phi_{aS}(u)$ holds in this case. This can also be observed by the corresponding new substructure in form of Figure 2 that appears in $\mathsf{STree}(aS)$. Clearly $(a,b)\notin\Phi_{S}(u)$ because $\mathsf{occ}_{S}(aub)=0$. Hence $u$ is of Type (i) when the aforementioned conditions are satisfied.
Next, we show that $u$ is the only possible node of Type (i). Let $u^{\prime}$ be any node on the path between $v$ and $u$, exclusive. Notice that these nodes $u^{\prime}$ are the only nodes other than $u$ that obtain a new left-extension $a$ (see also Figure 1). Assume for a contrary that $u^{\prime}$ is of Type (i), which implies that $\mathsf{occ}_{aS}(au^{\prime})=1$ and $\mathsf{occ}_{S}(au^{\prime})=0$. Let $b^{\prime}$ be the only character that immediately follows $au^{\prime}$ in $aS$, namely $\mathsf{occ}_{aS}(au^{\prime}b^{\prime})=1$. Note that such character $b^{\prime}$ must exist for each $u^{\prime}$. It follows from Lemma 6 that it suffices to check whether $u^{\prime}$ has an out-edge labeled $b^{\prime}$ that leads to a leaf. However, this is not the case: $au^{\prime}b^{\prime}$ is a prefix of $aS$, and thus $u^{\prime}b^{\prime}$ is a prefix of $S$. Thus, $b^{\prime}$ is the first character in the path label from $u^{\prime}$ to the leaf $S$. Since $u^{\prime}$ is a non-parent ancestor of the leaf $S$, the out-edge from $u^{\prime}$ that begins with $b^{\prime}$ does not lead to a leaf. Thus, $u^{\prime}$ cannot be a Type (i) node by Lemma 6. ∎
This node $u$ is found in the process of climbing up the path from the leaf for $S$ toward its lowest ancestor $z$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ for $a\in\Sigma$. The character $b$ can easily be obtained by using the string depths of $u$ and $S$. When all the conditions in Lemma 7 are satisfied, then $u$ is of Type (i). We then update $\Phi_{aS}(u)\leftarrow\Phi_{S}(u)\cup\{(a,b)\}$. We add $(u,1)$ to the list $\mathcal{L}$ if $\phi_{S}(u)=0$, and replace $(u,\phi_{S}(u))$ with $(u,\phi_{S}(u)+1)=(u,\phi_{aS}(u))$ in $\mathcal{L}$ otherwise.
@@ 3.5 Dealing with Type (ii) node
Figure 3: A snapshot of Weiner’s suffix tree construction algorithm in the case where $ax$ is a leaf. If the leaf $x$ is a child of $v$, then $(a,c)\in\Phi_{S}(v)\setminus\Phi_{aS}(v)$. Let $d$ be a unique character that immediately precedes $ax$ in $S$. Then $(d,c)\in\Phi_{aS}(av)\setminus\Phi_{S}(av)$. If dazdaz in an implicit node on the edge leading to the leaf daxdax, then $(d,e)\in\Phi_{S}(az)\setminus\Phi_{aS}(az)$.
@@
Recall that $av$ becomes an internal node as the parent of a new leaf for $aS$ in $\mathsf{STree}(aS)$ after the update. Since Weiner’s algorithm inserts at most one internal node and it is $av$ in this case, the string davdav remains on an edge in $\mathsf{STree}(aS)$ after the update. Thus, $av$ obtains a soft Weiner link labeled $d\in\Sigma$ toward node daxdax. Since $ax$ is a leaf, daxdax is also a leaf. For the string $t$ such that $dax=davt$, we have $t[1]=c$. Thus, we have $(d,c)\in\Phi_{aS}(av)$ by Lemma 6. This can also be observed by the corresponding new substructure in form of Figure 2 that appears in $\mathsf{STree}(aS)$. Due to Lemma 3, $d$ is the unique character pair for $c$ that can possibly be a member of $\Phi_{aS}(av)$. Since $c$ was the only character immediately following $av$ in $S$, $(d,c)$ is the only character pair that could be a member for $\Phi_{aS}(av)$. Since $\mathsf{occ}_{S}(av)=1$, $av$ was not a repeat in $S$ before the update. Thus $\phi_{S}(av)=0$ and hence $(d,c)\notin\Phi_{S}(av)$ holds for the string $S$ before the update.
If $ax$ is not a leaf, then $\mathsf{occ}_{aS}(avc)\geq\mathsf{occ}_{aS}(ax)>1$, implying $(d,c)\notin\Phi_{aS}(av)$. Also, if $av$ is a node in $\mathsf{STree}(S)$, then $av$ was already a maximal repeat in $S$ and thus $av$ is not a Type (ii) node by definition. ∎
The locus for $av$ is found via the Weiner link ${\mathsf{W\_link}}_{a}({v})$ from the node $v$ during the update from $\mathsf{STree}(S)$ to $\mathsf{STree}(aS)$. We remark that $av$ is on an edge in $\mathsf{STree}(S)$ if and only if ${\mathsf{W\_link}}_{a}({v})$ is a soft Weiner link in $\mathsf{STree}(S)$ before the update. Further, we can easily check if $ax$ is a leaf or not. The character $c$ can be obtained by using the string depths of the nodes $av$ and $ax$ (see also Figure 3). The character $d$ can be found easily (e.g. by using the unique Weiner link from leaf $ax$.)
@@ 3.6 Dealing with Type (iii)-1 node
[…]
Next, we show that $v$ is the only possible node that is of Type (iii)-1. Let $v^{\prime}$ be any node in the path between $v$ and $z$ exclusive, where $z$ is the lowest ancestor of the leaf $S$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ labeled $a$ in $\mathsf{STree}(S)$ before the update. Note that those nodes $v^{\prime}$ are the only nodes other than $v$ in the path from the root to $S$, which satisfy $\mathsf{occ}_{S}(av^{\prime})=1$ and $\mathsf{occ}_{S}(v^{\prime})>1$. However, since the descendant $v$ of $v^{\prime}$ is an internal node, we have $\mathsf{occ}_{S}(v^{\prime}c^{\prime})>1$ and thus $(a,c^{\prime})\notin\Phi_{aS}(v^{\prime})$ holds, where $c^{\prime}$ is the only possible paring character for character $a$ with node $v^{\prime}$ due to Lemma 3. Thus $v$ is the only possible node of Type (iii)-1. ∎
@@
Recall that $z$ is the lowest ancestor of the leaf $S$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})$ labeled $a$ in $\mathsf{STree}(S)$ before the update. If $v\neq z$, then $av$ is on an edge. We can easily check whether $x$ and $ax$ are leaves and $x$ is a child of $v$ in $\mathsf{STree}(S)$. Notice that the first condition $v\neq z$ is critical, since if $v=z$, then avav (=az=az) is an internal node in $\mathsf{STree}(S)$ and thus $\mathsf{occ}_{S}(av)>1$. Now we have two sub-cases:
@@ 3.7 Dealing with Type (iii)-2 node
Recall that $az$ is the parent of $ax$ in $\mathsf{STree}(S)$ before the update. Consider the case that $\mathsf{occ}_{S}(aze)=1$, where $ax$ is a leaf in $\mathsf{STree}(S)$. Also, consider the case that $\mathsf{occ}_{S}(daz)=1$, which implies $\mathsf{occ}_{S}(daze)=1$. Therefore, $(d,e)\in\Phi_{S}(az)$ by Lemma 6. Consider the case where $av$ becomes a branching node in $\mathsf{STree}(aS)$ after the update. Then, $av$ becomes a child of $az$, meaning that $\mathsf{occ}_{aS}(aze)>1$. Thus $(d,e)\notin\Phi_{aS}(az)$ holds for $az$. Overall, if the aforementioned conditions are satisfied, then $az$ is a node of Type (iii)-2. This can also be observed with the corresponding existing substructure in form of Figure 2 that disappears in $\mathsf{STree}(aS)$.
@@ 3.8 Complexities
By maintaining $\mathcal{L}$ as a doubly-linked list such that each element is associated to the corresponding node in the suffix tree, we obtain the following: